OpenStudy (anonymous):

Help Help help Determine whether y varies inversely with x. If so find the constant of variation. y = 7/x. Select one: a. Yes, k equals 7 b. No c. Yes, k equals -7 d. Yes, k equals 1/7 Determine whether y varies inversely with x. If so find the constant of variation. y = 7x. Select one: a. No b. Yes, k equals -7 c. Yes, k equals 7 d. Yes, k equals 1/7

OpenStudy (anonymous):

For the first one, A and for the second one A.I am not 100% certain, but those are what seem the most logical and correct to me.

OpenStudy (anonymous):

WAIT I GOT IT WRONG!!

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

1 is B and 2 is C

OpenStudy (anonymous):

are you sure ? >.<

OpenStudy (lncognlto):

I would say @agross was right the first time.

OpenStudy (anonymous):

yes, take y=1/x. say y=4, x would have to equal .25, and if y=16, x would have to =1/16

OpenStudy (anonymous):

So as y gets bigger, x has to get smaller.

OpenStudy (anonymous):

OK I will try that out.

OpenStudy (anonymous):

so yes, i was right the first time...SORRY!!!!! I just got the numbers mixed up in my head lol...sorry!!

OpenStudy (anonymous):

Yes so the correct answers are a. and a.